OpenStudy (anonymous):

write each as a summation notation 1) 2/4 +2/9 + 2/16 + 2/25 + 2/36 2) –2 + 4 – 6 + 8 – 10 + 12 3) 1 + 1/4 + 1/9 + 1/16 + 1/25 4) –3 + 6 – 9 + 12 – 15 5) 1 + 10 + 100 + 1000 + 10,000 6) 100 + 95 + 90 + 85 + 80

OpenStudy (sleepyhead314):

Usually it is a little daunting to have to help on multiple questions at once as a Helper :) so just try to ask these questions one at a time next time ^_^ for these sorts of questions, try to find a pattern like for the first one, we see that the numerator (top) is always 2 and that the denominator (bottom) seems to increase in squares (2)^2 = 4 (3)^2 = 9 (4)^2 = 16 (5)^2 = 25 (6)^2 = 36 seeing this 2, 3, 4, 5, 6 (increasing by one) pattern means that we have found the thing to replace for "n" so for the first question it would kinda be like \[\sum_{n=2}^{\infty}\frac{ 2 }{ n^2 }\]where the weird "n=2" thing means that it is increasing by 1 each time and the starts with n=2 do you get it?

OpenStudy (anonymous):

i get that one but how do you do number 2

OpenStudy (sleepyhead314):

number 2 has a (-1)^n in the numerator

OpenStudy (sleepyhead314):

it's called an Alternating Series

OpenStudy (sleepyhead314):

the "n" is in the exponent for that because then if n is odd the number becomes negative and if n is even, the number becomes positive the 1st number was negative then the 2nd number was positive etc

OpenStudy (sleepyhead314):

the thing making the numbers in question 2 to go in the 2, 4, 6, 8, pattern is 2(n) so 2n * (-1)^n